# Cross-positive linear maps, positive polynomials and sums of squares

### v1

@@ 5.3. Example
In this subsection we give an explicit example of a “proper” cross-positive map $\Phi:M_{3}(\mathbb{R})_{\textrm{sym}}\to M_{3}(\mathbb{R})_{\textrm{sym}}$ built off Algorithm 5.5. Let
@@
The corresponding linear map $\Phi:M_{3}(\mathbb{R})_{\textrm{sym}}\to M_{3}(\mathbb{R})_{\textrm{sym}}$ is as follows:
@@
The polynomial $p_{\Phi}$ is nonnegative on $V(I)$ but not a sum of squares modulo $I$. Equivalently, $\Phi$ is a proper cross-positive map. This example was produced using Algorithm 5.5 starting with the points
@@
Appendix A Proof of Theorem 4.5
In the later proofs we will use the following elementary lemma ([KMŠZ19, Lemma 2.2]).
Lemma A.1.
Let $p,q\in\mathbb{N}$ be natural numbers such that $p>q.$ For every $n\in\mathbb{N}$ we have
$\binom{pn}{qn}^{\frac{1}{qn}}<\frac{p}{q}\left(\frac{p}{p-q}\right)^{\frac{p-q}{q}}.$
We denote $\mathcal{K}=\widetilde{\operatorname{Pos}}^{(n)}_{\mathcal{Q}}$. In the same way as in [KMŠZ19, §2.1.1] we establish that
To prove this claim we will use [Bar02, Corollary 2]. Let $(\mathbb{R}^{n}\otimes\mathbb{R}^{n})^{\otimes 2}$ be the $2$-nd tensor power of $\mathbb{R}^{n}\otimes\mathbb{R}^{n}$ Let $e_{1},e_{2}\in\mathbb{R}^{n}$ be standard unit vectors and let $w$ be the tensor
The group $\operatorname{SO}(n)$ acts on $(\mathbb{R}^{n}\otimes\mathbb{R}^{n})^{\otimes 2}$ by the natural diagonal action, i.e., for $g\in\operatorname{SO}(n)$ and all $x_{i}\in\mathbb{R}^{n}$ we have
$g(x_{1}\otimes\cdots\otimes x_{4})=gx_{1}\otimes\cdots\otimes gx_{4}$
and extend by linearity. We also define
and we integrate w.r.t. the Haar measure $\widehat{\sigma}$ on $\operatorname{SO}(n)$. Similarly as in [BB05, Example 1.2] we proceed as follows:
  1. (1)
    Note also that $\operatorname{SO}(n)(e_{1}\otimes e_{2})=T$.
  2. (2)
  3. (3)
    We introduce an inner product on $V_{2}$ by defining
[…]
By [Bar02, Corollary 2],
where the equality follows as in [Bar02, p. 404]. If $n$ is odd, we let $2k_{0}=9(n-1)$. Otherwise take $2k_{0}=9n$ to get
$D_{k_{0}}^{\frac{1}{2k_{0}}}\leq\binom{\frac{20}{9}k_{0}}{2k_{0}}^{\frac{1}{k_{0}}}.$
Since $2k_{0}=9\ell_{0}$ for some $\ell_{0}\in\mathbb{N}$ we get
where we used Lemma A.1 in the last inequality.
(A.1) $A=\int_{S_{\mathcal{M}}}\left\|f\right\|_{2k_{0}}\;{\rm d}\widetilde{\mu}=\int_{S_{\mathcal{M}}}\left(\int_{T}f^{2k_{0}}\;{\rm d}\sigma\right)^{\frac{1}{2k_{0}}}{\rm d}\widetilde{\mu}.$
Notice that
(A.2) $\int_{S_{\mathcal{M}}}\left(\int_{T}f^{2k_{0}}\;{\rm d}\sigma\right)^{\frac{1}{2k_{0}}}{\rm d}\widetilde{\mu}=\int_{c\in S_{V^{\ast}_{2}}}\left(\int_{g\in\operatorname{SO}(n)}\langle c,gv\rangle^{2k_{0}}{\rm d}\widehat{\sigma}(g)\right)^{\frac{1}{2k_{0}}}{\rm d}\breve{\sigma}(c),$
where $S_{V^{\ast}_{2}}$ is the unit sphere in $V^{\ast}_{2}$ endowed with the rotation invariant probability measure $\breve{\sigma}$. Combining (A.1), (A.2) we obtain
where we used [BB05, Lemma 3.5] for the inequality and [BB05, Remark p. 62] for the last equality. This equality proves the Claim and establishes the lower bound in Theorem 4.5. $\blacksquare$
Appendix B Proof of Theorem 4.6
In the same way as in [KMŠZ19, §2.3.1] it follows that
$H_{f}(g)=\langle f,g^{2}\rangle\quad\text{for }g\in\mathcal{U}.$
Note that
$\|f\|_{\operatorname{sq}}=\|H_{f}\|_{\infty}.$
In the same way as in [KMŠZ19, p. 3343–3344] (for $k_{1}=k_{2}=1$) it follows that
where the last inequality follows by $\|g^{2}\|_{2}=\|g\|_{4}^{2}$ and Proposition 4.11. To prove the Claim it remains to establish
(B.1) $\sqrt{\frac{2D_{\mathcal{U}}}{D_{\mathcal{M}}}}\leq 2^{\frac{3}{2}}n^{-1}.$
The dimensions $D_{\mathcal{U}}$, $D_{\mathcal{M}}$ are easily verified to be
$D_{\mathcal{U}}$ $=$ $\dim\mathbb{R}[{\tt x},{\tt y}]_{1,1}-1=n^{2}-1,$
Observe that
$\underbrace{\leq}_{n>2}$ $\frac{2^{3}}{n^{2}+2n-2}\leq\frac{2^{3}}{n^{2}},$
which proves (B.1). $\blacksquare$

### v2

@@ 5.3. Example
In this section we give an explicit numerical example of a “proper” cross-positive map $\widetilde{\Phi}:M_{3}(\mathbb{R})\to M_{3}(\mathbb{R})$ built using our ad-hoc implementation of Algorithm 5.5 in Wolfram Mathematica. Working with rational random data as per Algorithm 5.5 quickly leads to very large denominators with bad conditioning, necessitating working with floating point numbers. Let
@@
The corresponding linear map $\Phi:\mathbb{S}_{3}(\mathbb{R})\to\mathbb{S}_{3}(\mathbb{R})$ is as follows:
@@
The polynomial $p_{\Phi}$ is nonnegative on $V(I)$ but not a sum of squares modulo $I$. Equivalently, an arbitrary *-linear extension $\widetilde{\Phi}:M_{3}(\mathbb{R})\to M_{3}(\mathbb{R})$ of $\Phi$ is a proper cross-positive map, e.g., $\widetilde{\Phi}$ is trivial on antisymmetric matrices. This example was produced using Algorithm 5.5 starting with the points
@@
Acknowledgments
The authors thank the anonymous referees and the editors for their careful reading and insightful comments that have greatly improved this manuscript.